OpenStudy (aravindg):

For a given value of k,the products the roots of [x ^{2}-2kx+3k ^{2}-4=0 is 5 then roots may be characterized as ...............

OpenStudy (anonymous):

\[x^2-2kx+3k^2-4=0\] us that what this says?

OpenStudy (aravindg):

yews

OpenStudy (aravindg):

plz help

OpenStudy (anonymous):

gimme a second

OpenStudy (aravindg):

u thr?

OpenStudy (anonymous):

yeah thinking.

OpenStudy (aravindg):

u hav gtalk?

OpenStudy (anonymous):

when it says "characterized" what kind of answer? real? repeated? complex? what is the topic?

OpenStudy (aravindg):

quadratics

OpenStudy (aravindg):

it means its property

OpenStudy (anonymous):

yeah i got that, i mean what kind of answer is expected?

OpenStudy (aravindg):

it means its property

OpenStudy (aravindg):

real or irrational like that

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

product of roots is the constant, so \[3k^2-4=5\] \[3k^2=9\] \[k^2=3\] \[k=\pm\sqrt{3}\]

OpenStudy (anonymous):

equation is \[x^2-2\sqrt{3}x+5=0\] or \[x^2+2\sqrt{3}+5=0\]

OpenStudy (aravindg):

so.......

OpenStudy (anonymous):

guess we don't have to characterize them, we can find them exactly.

OpenStudy (aravindg):

u hav gtalk??plz give email id

OpenStudy (anonymous):

there is chat here

OpenStudy (anonymous):

bottom right

OpenStudy (aravindg):

wel its slow

OpenStudy (aravindg):

i hav many questions

OpenStudy (anonymous):

actually pretty quick. lets find the zeros ready?

OpenStudy (aravindg):

is it real and distinct roots

OpenStudy (anonymous):

we have solved for k and know it is \[k=\sqrt{3}\] or \[k=\sqrt{3}\] so equation is \[x^2-2\sqrt{3}x+5=0\] or \[x^2+2\sqrt{3}x+5=0\] we solve the first one using formula get \[x=\frac{2\sqrt{3}\pm+\sqrt{(2\sqrt{3})^2-4\times 1\times 5}}{2}\] \[x=\frac{2\sqrt{3}\pm\sqrt{-8}}{2}\] \[x=\frac{2\sqrt{3}\pm2\sqrt{2}i}{2}\] \[x=\sqrt{3}\pm \sqrt{2}i\]

OpenStudy (anonymous):

those are the solutions roots exactly. they are \[\{\sqrt{3}+\sqrt{2}i,\sqrt{3}-\sqrt{2}i\}\]